sin(^2)x(tan(^2)x+1

Answer by stanbon(48564)

sin(^2)x(tan(^2)x+1
-----------
Since tan^2x+1= sec^2x you get:
sin^2x*sec^2x=(sin^2x)/(cos^2x)= tan^2x
Cheers,
Stan H.